# The Atacama Cosmology Telescope: The Development of Machine Learning Tools for Detecting Millimeter Sources in Timestream Pre-processing

### v1

@@ Abstract
We present a new machine learning algorithm for classifying short-duration features in raw time ordered data (TODs) of cosmic microwave background survey observations. The algorithm, specifically designed for the Atacama Cosmology Telescope (ACT), works in conjunction with the previous TOD preprocessing techniques that employ statistical thresholding to indiscriminately remove all large spikes in the data, whether they are due to noise features, cosmic rays, or true astrophysical sources in a process called “data cuts”. This has the undesirable effect of excising real astrophysical sources, including transients, from the data. The machine learning algorithm demonstrated in this work uses the output from these data cuts and is able to differentiate between electronic noise, cosmic rays, and point sources, enabling the removal of undesired signals while retaining true astrophysical signals during TOD pre-processing. We achieve an overall accuracy of 90% in categorizing data spikes of different origin and, importantly, 94% for identifying those caused by astrophysical sources. Our algorithm also measures the amplitude of any detected source seen more than once and produces a sub-minute to minute light curve, providing information on its short timescale variability. This automated algorithm for source detection and amplitude estimation will be particularly useful for upcoming surveys with large data volumes, such as the Simons Observatory.
@@ 1 Introduction
Ground-based microwave astronomy has experienced a revolution in the past two decades, thanks to the deployment of large aperture telescopes that have thousands of detectors, increased their resolution, and advances in detector and read-out technology that have drastically improved their sensitivity while also allowing for simultaneous observation of multiple wavelength bands. For example, the Atacama Cosmology Telescope (ACT; Swetz et al. 43, Thornton et al. 45, Henderson et al. 25) observed the cosmic microwave background (CMB) with 3000 detectors across up to three frequency bands, covering six bands from 30 to 270 GHz during its operation from 2007 to 2022. ACT will be succeeded by the Simons Observatory (SO), which is initially set to deploy over 60,000 detectors split between large- and small-aperture telescopes across six frequency bands from 30 to 280 GHz 3, 2. During observations, detectors, such as transition edge sensors (TES), record the incident power at a cadence of a few hundred hertz as the telescope scans in azimuth at a constant elevation. In the ACT survey, the data timestreams of all detectors are packaged into files known as “time-ordered data streams” (TODs), each containing roughly 11 minutes worth of data. Using knowledge of the telescope pointing, which is recorded during data collection, the TODs are later processed into microwave sky maps using maximum likelihood map-making 19, 17, 4.
@@
In addition to using the CMB directly as a probe of cosmological physics, an emerging use of CMB experiments is the study of the time-varying microwave sky. Wide-area surveys contain thousands of radio sources; the majority of them are blazars that emit microwave light that changes in flux over days/weeks. The South Pole Telescope (SPT) collaboration 15 published a pilot study of one blazar (PKS 2326-502) observed in the microwave bands 28. Surveys like ACT and SO will provide much longer baselines of observations of such objects: the field of mm-transient studies is expected to grow significantly in the coming decade. Similarly in the realm of planetary science, ACT and SPT have also characterized the millimeter flux of asteroids 16, 39, and ACT constrained the existence of a “Planet 9” in our Solar System 36. Finally, ACT and SPT have successfully measured dozens of transients in their data that are mainly flaring stars, with the majority consisting of short-duration (minutes- to hours-long) stellar flares 49, 37, 31, 22, 44, 12. Additionally, as part of a blind transient search, ACT has detected classical nova YZ Ret during its outburst, making it only the second millimeter observation of this kind 12. ACT also published a targeted search for extragalactic transients that included gamma ray bursts (GRBs), tidal disruption events (TDEs), and supernovae (SN), placing upper-limits on the microwave flux of the target objects 26. More sensitive data from the upcoming SPT, SO, and CMB-S4 1 surveys will produce detections in the millimeter of extragalactic transients, specifically GRBs, which will provide valuable information on the emission processes of these powerful explosions 20, 2.
@@
CMB surveys are designed to measure the power of faint microwave radiation imprinted on all angular scales across the sky. However, the data collected by ground-based CMB telescope like ACT are dominated by microwave emission from water vapor in the atmosphere and are additionally contaminated from, e.g., instrumental noise and cosmic rays colliding with the detectors. While instrumental and atmospheric noise can be modeled in the microwave map-making process to produce unbiased, maximum likelihood maps 19, the mapmaker is specifically vulnerable to biases from short, high-amplitude, non-Gaussian data bursts in the TOD, and therefore depends on these bursts being excised from the TOD before map-making. Such transient events can be caused by sporadic pathologies from the electronic readout system, cosmic rays, electromagnetic interference, radio frequency interference modulated by telescope motion, ionizing radiation, digitization artifacts, etc 9, 8, 33. Cosmic rays, for example, sometimes produce rapid spikes in the TODs of single detectors and may sometimes also warm up a region of the detector polarization array (PA) near the impact, thereby inducing a signal spike in multiple detectors (see 33 for a study on cosmic rays for Planck HFI). In this paper, we refer to all of these transient phenomena as “glitches”. Bright astrophysical point sources can also cause high-amplitude, non-Gaussian spikes in the TOD and may be mistaken for a glitch. This may be avoided if the coordinate of the source is known before pre-processing. However, this also means that any point source or astrophysical transient that is not already known in advance of map-making can be removed from the data set entirely. We note that while the specific phenomena we describe here are commonly encountered in ACT timestreams, other transient sources may manifest differently in different telescopes depending on their readout schemes and observing conditions.
@@
It is therefore imperative when pre-processing TODs from CMB telescopes to excise as many glitches as possible while still retaining as many true astrophysical transients as feasible. As the volume of data observed by current and future experiments increases, traditional techniques that rely on human inspection of a significant fraction of candidate glitches in order to tune statistically-based thresholding algorithms will rapidly become unwieldy. Improvements in the data-excision algorithms are needed, both to better automate the process and to avoid cutting real signals. In this paper, we describe a new machine learning method that achieves these objectives utilizing the outputs of the current glitch finder.
@@
After reviewing in Section 2 the “cuts process” used by ACT to cut out potential glitches from the data, we describe our new approach in Sections 3 and 4, and introduce our supervised machine learning algorithm in Section 4. We explain the summary statistics derived from the TODs that are fed into the algorithm in Section 4 and how we prepare our training and test sets in Section 5. We test the classification method on simulated stellar flares which are discussed in Section 6, summarize and discuss our results in Section 7, and present our conclusions and suggestions for future work in Section 8.
@@ 2 Current State of ACT Data Cuts
Firstly, the data cuts algorithm can discard entire TODs based on the level of precipitable water vapor in the atmosphere during the TOD, or the number of faulty detectors operating over that time period. Additionally, we flag individual detectors within each TOD based on various statistical metrics computed on the TOD. A detector may become faulty for various reasons, such as saturation from optical and electric power, an instability caused by insufficient heat capacity, slow time constants (which describes how quickly the detector changes in response to a signal) from fabrication and wiring problems, or failures in the detector readout system. Current methods for producing individual detector cuts utilize a set of parameters to identify and exclude such detectors with anomalous behavior. These anomalies often manifest as abnormally high or low noise levels, poor optical coupling, or excessive high-frequency, non-Gaussian noise above around 10 Hz, where photon shot noise with a white spectrum is expected to dominate. Detectors that have outlier values of the statistics for these parameters are flagged and rejected from map-making.
@@
Moreover, short, high signal-to-noise (SNR) spikes on millisecond (ms) time scales are also prevalent in the detector timestreams, caused by glitches in the readout system and cosmic rays, among other factors. ACT employs a glitch finder that combines high-pass and Gaussian filters to identify samples affected by glitches, defined as those with in the high-pass-filtered TOD. This method is effective because ACT TODs are predominantly noise-dominated. When a glitch is flagged, a buffer of 200 time samples on either end of the glitch is identified, and all samples from the detector within this buffer interval are rejected from the map-making process.
@@
While this glitch finder is effective at removing glitches due to non-astrophysical signals, it also excludes bright astronomical sources that appear as spikes in the timestream as the telescope scans across them. Given the ACT scanning speed () and beam size arcminute at 150 GHz, astrophysical source spikes typically last around 20–30 ms. In the standard map-making pipeline, a catalog of bright astrophysical sources is curated and passed to the glitch finder; samples in the TOD that lie within of a known source from the catalog are masked before running the map-maker. This approach ensures that known bright astrophysical sources will not be cut, but it necessitates multi-pass TOD processing and, more importantly, may fail when a previously unknown source of transient nature appears in the observing field, such as those reported in Naess et al. 37 and Li et al. 31. If these signals are bright enough, they might be erroneously removed by the glitch finder, resulting in missed opportunities to discover new sources of astrophysical interest.
@@
Different types of glitches typically manifest differently in the timestream and across the focal plane. For instance, cosmic rays often simultaneously affect detectors in close proximity to each other. In contrast, a signal from a bright astrophysical point source is extended on the focal plane given the finite beam-shape. Point source signals are thus extended in the timestream as well and appear at different times for different detectors depending on their locations on the focal plane. As the telescope scans the sky at a constant elevation, detectors that see the same source fall along a constant-elevation line on the focal plane during each scan. The distinct spatial and temporal characteristics of different glitches can in principle be used to differentiate them. The spatio-temporal information of the TOD data has also been successfully used to model the bulk motion of the atmosphere 34.
@@
In this work, we demonstrate that by combining both spatial and temporal information for different classes of glitches, it is possible to effectively distinguish between them with the assistance of a novel machine learning model. This approach provides insights into the physical origins of each glitch, and offers a promising method for preserving and identifying astrophysically interesting transient signals in the TODs in a fully automated method without the need for visual inspection by human experts. This property of the glitch classification will have growing importance as the data volume scales up in the coming decade with surveys like SO.
@@ 3 Data Preparation
Figure 1 shows snippets for the four categories we define in this work to generally summarize the phenomena we see in the glitches for classification: point sources (PS), point sources with another coincident glitch (abbreviated as PS+), cosmic rays (CR), and electronic glitches (EG). In the figure, the example snippet for the PS+ contains a point source together with a cosmic ray, but this category can also include an electronic glitch rather than a cosmic ray. There is a large variety of electronic glitches that are observed in our data (all labeled as EG): for instance, in the bottom panel of Figure 1, a diagonal “striping” is seen on the focal plane. These detectors share a common read-out line in the time-domain multiplexed system of our Multi-Channel Electronics (MCE) readout system 10. This particular electronic glitch has affected the whole read-out line. For our machine learning classification of these categories, we characterize both the TODs and layout of the affected detectors across the focal plane using the summary statistics described in the following section.
4 Description of Classification Features and Hyperparameters
@@
the number of detectors affected by the glitch (Number of Detectors; Section 4.1);
@@
summaries of the distribution of these detectors across the focal plane (Y and X Extent Ratio, Normalized Y Maximum, and Normalized Y Maximum Within 0.1; Section 4.2);
@@
the correlation of the signal (Correlation; Section 4.3);
@@
the time lag between detectors (Time-lag; Section 4.4);
@@
and the number of peaks in the TODs (Number of Peaks; Section 4.5).
Figure 2 shows the distributions of the seven statistics, described in the subsections that follow, falling in each of the five types enumerated above.
@@ 4.1 Number of Detectors
We consider the number of detectors affected by a glitch as one of our summary statistics. Generally for a point source we expect (but generally fewer than 100) detectors to be affected depending on where the point source crosses the focal plane as the telescope scans in azimuth, as only a strip of detectors will observe the source. A cosmic ray will generally only affect a small cluster of detectors; but they can affect large portions of the focal plane, depending on the energy of the cosmic ray and if multiple cosmic rays hit the focal plane at the same time. The largest spread in the number of affected detectors occurs for the EG category, as it includes a large variety of glitches. Many electronic glitches will only affect detectors, however there are instances, for example large-area heat-ups, where almost every detector on the focal plane is affected. Finally, when a point source occurs coincident with another glitch, we expect a slightly larger number of affected detectors than for the PS category, since there is another glitch occurring as well.
@@ 4.2.1 The and Extent Ratio
Since many detectors will observe a PS across the focal plane in a horizontal line as the telescope scans in azimuth, we expect a large spread in the coordinate but only a small range of affected detectors in the direction, since the celestial sphere barely rotates during a single scan. Therefore, there should generally be a small ratio of extent compared to the extent/range for a point source. A cosmic ray will hit a cluster of detectors, which results in a similar spread in the and directions, and we therefore generally expect the ratio to have a value close to unity for CR objects. However, this is not always the case. For example, if multiple cosmic rays hit the focal plane at once, the ratio will deviate from unity, but the “ and Extent Ratio” will typically remain larger than that of a point source. The expected extent ratio also varies quite significantly for the EG category depending on the type of electronic glitch, however the ratio is generally larger than that for a point source. For a PS+, we expect a variety of deviations that differ from the extent ratio expected for a point source alone, depending on the location of the other glitch with respect to the location of the source on the focal plane. If the glitch occurs near the source, we similarly expect a small ratio. Conversely, if the glitch occurs far away on the focal plane, the ratio will be closer to one as it will have a large range and the range is generally close to the full extent of the focal plane for a point source.

### v2

# The Atacama Cosmology Telescope: Machine Learning Driven Tools for Detecting Millimeter Sources in Timestream Pre-processing

@@ Abstract
We present a new pipeline utilizing machine learning for classifying short-duration features in raw time-ordered data (TOD) of cosmic microwave background survey observations. The pipeline, specifically designed for the Atacama Cosmology Telescope, works in conjunction with the previous TOD preprocessing techniques that employ statistical thresholding to indiscriminately remove all large spikes in the data, whether they are due to noise features, cosmic rays, or true astrophysical sources, in a process called “data cuts”. This has the undesirable effect of excising real astrophysical sources, including transients, from the data. The classification pipeline demonstrated in this work uses the output from these data cuts and is able to differentiate between electronic noise, cosmic rays, and point sources, enabling the removal of undesired signals while retaining true astrophysical signals during TOD preprocessing. We achieve an overall accuracy of 90% in categorizing data spikes of different origin and, importantly, 94% for identifying those caused by astrophysical sources. Our pipeline also measures the amplitude of any detected source seen more than once and produces a subminute-to-minute light curve, providing information on its short timescale variability. This automated pipeline for source detection and amplitude estimation will be particularly useful for upcoming surveys with large data volumes, such as the Simons Observatory.
@@ 1 Introduction
Ground-based microwave astronomy has experienced a revolution in the past two decades, thanks to the deployment of large aperture telescopes that have thousands of detectors, increased their resolution, and advances in detector and read-out technology that have drastically improved their sensitivity while also allowing for simultaneous observation of multiple wavelength bands. For example, the Atacama Cosmology Telescope (ACT; Swetz et al. 42, Thornton et al. 45, Henderson et al. 25) observed the cosmic microwave background (CMB) with 3000 detectors across up to three frequency bands, covering six bands from 30 to 270 GHz during its operation from 2007 to 2022.111Jupyter notebooks outlining how to use the latest ACT data products (ACT DR6) can be seen at https://github.com/ACTCollaboration/DR6_Notebooks, including a glitch classification notebook outlining the methods discussed in this paper. ACT will be succeeded by the Simons Observatory (SO), which is initially set to deploy over 60,000 detectors split between large- and small-aperture telescopes across six frequency bands from 30 to 280 GHz 3, 2. During observations, detectors, such as transition edge sensors, record the incident power at a cadence of a few hundred hertz as the telescope scans in azimuth at a constant elevation. In the ACT survey, the data timestreams of all detectors are packaged into files known as “time-ordered data” (TOD), each containing roughly 11 minutes worth of data. Using knowledge of the telescope pointing, which is recorded during data collection, the TODs are later processed into microwave sky maps using maximum likelihood map-making 19, 17, 4.
@@
In addition to using the CMB directly as a probe of cosmological physics, an emerging use of CMB experiments is the study of the time-varying microwave sky. Wide-area surveys contain thousands of radio sources; the majority of them are blazars that emit microwave light that changes in flux over days/weeks. The South Pole Telescope (SPT) collaboration 15 published a pilot study of one blazar (PKS 2326-502) observed in the microwave bands 28. Surveys like SO will provide much longer baselines of observations of such objects: the field of mm-transient studies is expected to grow significantly in the coming decade. Similarly in the realm of planetary science, ACT and SPT have also characterized the millimeter flux of asteroids 16, 39, and ACT constrained the existence of a “Planet 9” in our Solar System 36. Finally, ACT and SPT have successfully measured dozens of transients in their data that are mainly flaring stars, with the majority consisting of short-duration (minutes- to hours-long) stellar flares 48, 37, 31, 22, 43, 12. Additionally, as part of a blind transient search, ACT has detected classical nova YZ Ret during its outburst, making it only the second millimeter observation of this kind 12. ACT also published a targeted search for extragalactic transients that included gamma ray bursts (GRBs), tidal disruption events (TDEs), and supernovae (SN), placing upper-limits on the microwave flux of the target objects 26. More sensitive data from the upcoming SPT, SO, and CMB-S4 1 surveys will produce detections in the millimeter of extragalactic transients, specifically GRBs, which will provide valuable information on the emission processes of these powerful explosions 20, 2.
@@
CMB surveys are designed to measure the power of faint microwave radiation imprinted on all angular scales across the sky. However, the data collected by ground-based CMB telescope like ACT are dominated by microwave emission from water vapor in the atmosphere and are additionally contaminated from, e.g., instrumental noise and cosmic rays colliding with the detectors. While instrumental and atmospheric noise can be modeled in the microwave map-making process to produce unbiased, maximum likelihood maps 19, the mapmaker is specifically vulnerable to biases from short, high-amplitude, non-Gaussian data bursts in the TOD, and therefore depends on these bursts being excised from the TOD before map-making. Such transient events can be caused by sporadic pathologies from the electronic readout system, cosmic rays, electromagnetic interference, radio frequency interference modulated by telescope motion, ionizing radiation, digitization artifacts, etc 34, 8, 9. Cosmic rays, for example, sometimes produce rapid spikes in the TODs of single detectors and may sometimes also warm up a region of the detector polarization array (PA) near the impact, thereby inducing a signal spike in multiple detectors (see 34 for a study on cosmic rays for Planck High Frequency Instrument). In this paper, we refer to all of these transient phenomena as “glitches”. Bright astrophysical point sources can also cause high-amplitude, non-Gaussian spikes in the TOD and may be mistaken for a glitch. This may be avoided if the coordinate of the source is known before pre-processing. However, this also means that any point source or astrophysical transient that is not already known in advance of map-making can be removed from the data set entirely. We note that while the specific phenomena we describe here are commonly encountered in ACT timestreams, other transient sources may manifest differently in different telescopes depending on their readout schemes and observing conditions.
@@
It is therefore imperative when preprocessing TODs from CMB telescopes to excise as many glitches as possible while still retaining as many true astrophysical transients as feasible. As the volume of data observed by current and future experiments increases, traditional techniques that rely on human inspection of a significant fraction of candidate glitches in order to tune statistically based thresholding algorithms will rapidly become unwieldy. Improvements in the data-excision algorithms are needed, both to better automate the process and to avoid cutting real signals. In this paper, we describe a new method utilizing machine learning that achieves these objectives utilizing the outputs of the current glitch finder.
@@
After reviewing in Section 2 the “cuts process” used by ACT to cut out potential glitches from the data, we describe our data preparation in Section 3 and introduce the supervised machine learning algorithm in Section 4. We explain the summary statistics derived from the TODs that are fed into the classifier in Section 4 and how we prepare our training and test sets in Section 5. We test the classification method on simulated stellar flares which are discussed in Section 6, summarize and discuss our results in Section 7, and present our conclusions and suggestions for future work in Section 8.
@@ 2 Current State of ACT Data Cuts
First, the data cuts algorithm can discard entire TODs based on the level of precipitable water vapor in the atmosphere during the TOD, or the number of faulty detectors operating over that time period. Additionally, we flag individual detectors within each TOD based on various statistical metrics computed on the TOD. A detector may become faulty for various reasons, such as saturation from optical and electric power, an instability caused by insufficient heat capacity, slow time constants (which describes how quickly the detector changes in response to a signal) from fabrication and wiring problems, or failures in the detector readout system. Current methods for producing individual detector cuts utilize a set of parameters to identify and exclude such detectors with anomalous behavior. These anomalies often manifest as abnormally high or low noise levels, poor optical coupling, or excessive high-frequency, non-Gaussian noise above around 10 Hz, where photon shot noise with a white spectrum is expected to dominate. Detectors that have outlier values of the statistics for these parameters are flagged and rejected from map-making.
@@
Moreover, short, high signal-to-noise (SNR) spikes on millisecond (ms) time scales are also prevalent in the detector timestreams, caused by glitches in the readout system and cosmic rays, among other factors. ACT employs a glitch finder that combines high-pass and Gaussian filters to identify samples affected by glitches, defined as those with in the high-pass-filtered TOD. This method is effective because ACT TODs are predominantly noise dominated. When a glitch is flagged, a buffer of 200 time samples on either end of the glitch is identified, and all samples from the detector within this buffer interval are rejected from the map-making process.
@@
While this glitch finder is effective at removing glitches due to non-astrophysical signals, it also excludes bright astronomical sources that appear as spikes in the timestream as the telescope scans across them. Given the ACT scanning speed () and beam size at 150 GHz, astrophysical source spikes typically last around 20–30 ms. In the standard map-making pipeline, a catalog of bright astrophysical sources is curated and passed to the glitch finder; samples in the TOD that lie within of a known source from the catalog are masked before running the map-maker. This approach ensures that known bright astrophysical sources will not be cut, but it necessitates multi-pass TOD processing and, more importantly, may fail when a previously unknown source of transient nature appears in the observing field, such as those reported in Naess et al. 37 and Li et al. 31. If these signals are bright enough, they might be erroneously removed by the glitch finder, resulting in missed opportunities to discover new sources of astrophysical interest.
@@
Different types of glitches typically manifest differently in the timestream and across the focal plane. For instance, cosmic rays often simultaneously affect detectors in close proximity to each other. In contrast, a signal from a bright astrophysical point source is extended on the focal plane given the finite beam-shape. Point source signals are thus extended in the timestream as well and appear at different times for different detectors depending on their locations on the focal plane. As the telescope scans the sky at a constant elevation, detectors that see the same source fall along a constant-elevation line on the focal plane during each scan. The distinct spatial and temporal characteristics of different glitches can in principle be used to differentiate them. The spatio-temporal information of the TOD data has also been successfully used to model the bulk motion of the atmosphere 35.
@@
In this work, we demonstrate that by combining both spatial and temporal information for different classes of glitches, it is possible to effectively distinguish between them with the assistance of machine learning. This approach provides insights into the physical origins of each glitch, and offers a promising method for preserving and identifying astrophysically interesting transient signals in the TODs in a fully automated method without the need for visual inspection by human experts. This property of the glitch classification will have growing importance as the data volume scales up in the coming decade with surveys like SO.
@@ 3 Data Preparation
Figure 1 shows snippets for the four categories we define in this work to generally summarize the phenomena we see in the glitches for classification: point sources (PS), point sources with another coincident glitch (abbreviated as PS+), cosmic rays (CR), and electronic glitches (EG). In the figure, the example snippet for the PS+ contains a point source together with a cosmic ray. There is a large variety of electronic glitches that are observed in our data (all labeled as EG): for instance, in the bottom panel of Figure 1, a diagonal “striping” is seen on the focal plane. These detectors share a common read-out line in the time-domain multiplexed system of our multi-channel electronics read-out system 10. We characterize both the TODs and layout of the affected detectors across the focal plane of the snippets using the summary statistics described in the following section and use them as the features for classification.
4 Description of Classification Features
@@
The number of detectors affected by the glitch (Section 4.1).
@@
Summaries of the distribution of these detectors across the focal plane ( and extent ratio, normalized maximum, and normalized maximum within 0.1; Section 4.2).
@@
The correlation of the signal (Section 4.3).
@@
The time lag between detectors (Section 4.4).
@@
The number of peaks in the TODs (Section 4.5).
Figure 2 shows the distributions of the seven statistics, described in the subsections that follow, falling in each of the five types enumerated above. Alternative unsupervised methods, that utilize focal plane images or the entirety of the temporal data for example, could be explored in future work. However, the goal here is to develop a computationally light method that could run on-site for future telescopes for real-time transient detection.
@@ 4.2.1 The and Extent Ratio
Since many detectors will observe a PS across the focal plane in a horizontal line as the telescope scans in azimuth, we expect a large spread in the coordinate but only a small range of affected detectors in the direction, since the celestial sphere barely rotates during a single scan. Therefore, there should generally be a small ratio of extent compared to the extent/range for a point source. A cosmic ray will hit a cluster of detectors, which results in a similar spread in the and directions, and we therefore generally expect the ratio to have a value close to unity for CR objects. However, this is not always the case. For example, if multiple cosmic rays hit the focal plane at once, the ratio will deviate from unity, but the and extent ratio will typically remain larger than that of a point source. The expected extent ratio also varies quite significantly for the EG category depending on the type of electronic glitch, however the ratio is generally larger than that for a point source. For a PS+, we expect a variety of deviations that differ from the extent ratio expected for a point source alone, depending on the location of the other glitch with respect to the location of the source on the focal plane. If the glitch occurs near the source, we similarly expect a small ratio. Conversely, if the glitch occurs far away on the focal plane, the ratio will be closer to one as it will have a large range and the range is generally close to the full extent of the focal plane for a point source.